# Exploiting Student Parallelism for Low-latency GPU Inference of BERT-like Models in Online Services

### v1

@@ Abstract.
Due to high accuracy, BERT-like models have been widely adopted by discriminative text mining and web searching. However, large BERT-like models suffer from inefficient online inference, as they face the following two problems on GPUs. First, they rely on the large model depth to achieve high accuracy, which linearly increases the sequential computation on GPUs. Second, stochastic and dynamic online workloads cause extra costs.
@@ 1. Introduction
Due to the marvelous success in natural language understanding, BERT-like models (devlin2018bert, ; liu2019roberta, ; he2021deberta, ) have been the foundation model for discriminative text mining and web searching in various online services. However, large BERT-like models pay the price of high inference latency for the model capability. Besides high accuracy, a wide range of web services are time-sensitive and require efficient online inference. For example, news sentiment classification for automatic stock trading should be quicker than any others, because even a sub-millisecond delay can determine the profits (aquilina2022quantifying, ); web searching usually has a ten-millisecond-level time budget for every query to score and rank the relevance of document candidates (yang2020model, ); and online writing assistants need to give immediate response to text streams while users are typing (DBLP:conf/acl/KanekoMKSI20, ).
@@
Some existing works focus on reducing the computation complexity of BERT-like models such as the compact model design (sun2020mobilebert, ; xu2021bert, ), knowledge distillation (sanh2019distilbert, ; sun2019patient, ; jiao2019tinybert, ), and pruning (DBLP:conf/rep4nlp/GordonDA20, ; michel2019sixteen, ; hou2020dynabert, ) which effectively reduce the model size for weak edge devices. However, since deep learning essentially relies on model depth for high model capability (lecun2015deep, ), these methods prefer reducing model width to model depth (about half the layers remained) to better maintain the accuracy. Moreover, early exiting methods (DBLP:conf/acl/XinTLYL20, ; DBLP:conf/acl/LiuZWZDJ20, ; zhou2020bert, ) still have some hard samples to go through all layers. Therefore, their deep architectures largely degrade their inference performance on GPUs which are the de-facto platform for deep learning models. Figure 1(a) shows that model depth linearly increases the latency due to serial forward propagation, while GPU has enough parallel computation capability to keep the same latency and throughput for different model widths as shown in Figure 1(b). Therefore, these deep and thin models are not suitable for efficient GPU inference.
@@
Furthermore, previous works pay extra costs for the stochastic online workloads. Different from collected data batches in training, online inference samples of different lengths arrive in stochastic and dynamic patterns. Since there are occasional workload bursts due to breaking events, previous works have to reserve over-provisioned GPUs idle for the most time (zhang2019mark, ; DBLP:conf/usenix/Romero0YK21, ; 276950, ). Additionally, Figure 1(c) indicates that GPUs need large batches for high throughput, because latency remains stable and throughput keeps increasing until the batch size is large enough. For high throughput and utilization, existing works employ a waiting queue to batch enough data samples from stochastic online workloads (zhang2019mark, ; ali2020batch, ; DBLP:conf/usenix/Cui00WLZ0G22, ), so early-arrived samples have higher latency due to the waiting. And text sequences have different lengths ranging from 10 to hundreds, but data parallelism on GPUs requires all samples in one batch to have the regular length. Existing works commonly pad short samples into the maximum length to form the regular batch (devlin2018bert, ). As shown in Figure 1(d), latency can increase rapidly with sequence length due to $O(N^{2})$ complexity of self-attention (DBLP:conf/nips/ZaheerGDAAOPRWY20, ), so paddings also bring in significant extra computations.
@@
Figure 1. Influences of different model and input factors on latency (red solid line) and throughput (green dashed line)
Therefore, we present Student Parallelism to explore student parallelism for low-latency GPU inference of BERT-like models, especially for the web services having stochastic online workloads. At the core of Student Parallelism, it distills the original deep model into a parallel group of shallow student models. Therefore, the sequential layer computations can be replaced with the equivalent parallel computations of students without compromising accuracy. To achieve this, we propose a novel knowledge distillation method based on explicit boosting ensemble and stacking distillation (§ 3.2). Despite the parallel inference, every student is sequentially trained to gradually reduce the error between the teacher and the group of students. Furthermore, every student distills all previous students into its intermediate layer by stacking distillation, so that all students can make better corrections with the estimated output of previous students. Therefore, all students can maintain the accuracy with no more than one quarter of the original layers (even only two layers in some cases), which significantly reduces the inference latency on GPUs. Additionally, it conducts adaptive student pruning to improve the generalization of the remained students and drop some last-added students with minimal accuracy loss, which can temporarily improve throughput for the occasional workload burst.
To reduce the extra costs of batching from online workloads, Student Parallelism employs specialized system designs for student parallelism (§ 3.3). It allocates the students of the same group on different GPUs to better parallelize even the single sample. It further replicates the student group into multiple copies to run concurrently and share the same set of GPUs with MPS. Therefore, it enables immediate and concurrent inference of different samples without batching and interference to minimize waiting and paddings.
@@
In summary, our work makes the following contributions:
@@
    We propose student parallelism to distill the deep model into a group of parallel students with lower depth than baselines (e.g., two layers) for low latency with comparable accuracy.
@@
    To handle online workloads, we can reduce the student number with minimal accuracy loss for workload burst, and we also make specialized system designs for student parallelism to reduce waiting and paddings.
@@
    We conduct comprehensive experiments to verify the effectiveness of our Student Parallelism, where it outperforms other baselines in terms of accuracy, latency, and throughput.
@@ 2. Related Work
  • Model Compression: these works only generate a single small model for weak edge devices, which still suffer from considerable model depth and high latency on GPUs. MobileBERT (sun2020mobilebert, ) and NAS-BERT (xu2021bert, ) make compact model design that can achieve the similar accuracy. Knowledge distillation transfers the knowledge from a large teacher model to a small student model (sanh2019distilbert, ; sun2019patient, ; jiao2019tinybert, ), so that the small student mimics the behavior of its teacher. Pruning can remove redundant weights, neurons, attention heads or tokens in well-trained models for less costs and better generalization (hou2020dynabert, ; malach2020proving, ; kim2022learned, ). Early exiting (DBLP:conf/acl/XinTLYL20, ; DBLP:conf/acl/LiuZWZDJ20, ; zhou2020bert, ) adds multiple classifiers on different layers as the multiple exits, so that it allows early exit from bottom layers for ”easy” samples. Quantization (kim2021bert, ; shen2020q, ; zafrir2019q8bert, ; tang2022mkq, ) replaces slow float operators with fast integer operators, which releases higher parallel computation capability. However, deep learning essentially relies more on model depth than model width for high model capability (lecun2015deep, ), so they still require a considerable large model depth to maintain the accuracy. For example, MobileBERT and NAS-BERT are still deep (12 or even 24 layers) but much thinner to avoid accuracy loss. 6-layer DistillBERT can maintain 96.3% overall score of the 12-layer Bert teacher, but the 3-layer PKD-BERT can only get 88.5% (§ 4.2.1). Model pruning (hou2020dynabert, ) can only reduce about the half of layers at most to maintain good accuracy. In early exiting, the ”difficult” samples still have to go through all layers. Quantization not only needs hardware support in low-bit integers (zafrir2019q8bert, ) but also fails to reduce the model depth and computation complexity. It also relies on large batch sizes to increase parallelism degree and speed-up (e.g., the speedup is 2.4×2.4$\times$2.4 × and $\times3$ when the batch size is 1 and 8 respectively (kim2021bert, )).
@@
  • Weak Model Ensemble: Clipper (DBLP:conf/nsdi/CrankshawWZFGS17, ) and Cocktail (276950, ) adopt bagging ensemble of different relatively small models to run in parallel for lower latency and high accuracy, which can only select the models readily trained in the above-mentioned methods. However, they rely on the different model architectures for the ensemble diversity, resulting in an imbalanced workload and straggler problem. The Mixture-of-Experts (MoE) in attention heads (DBLP:conf/iclr/ShazeerMMDLHD17, ; DBLP:conf/naacl/ZuoZLHZC22, ) has a router to activate only one or two experts in every layer. It can keep constant inference cost no matter how wide the model is, but it cannot decrease the model depth. SenseAI (wang2021sensai, ) divides the multi-class model into several pruned binary models, which still have a large model depth and irregular pruned architectures. It cannot be used in regression and binary classification tasks either. Some collaborative and self-distillation (DBLP:conf/icann/ZainJZ22, ; sun2021collaborative, ) explore distilling multiple teachers and existing students into a new better student for the time-shifting distribution, but they do not leverage multiple parallel students together for lower model depth.
@@
  • System Designs for Online Workloads: existing system works pay extra costs to handle dynamic online workloads. They reserve some over-provisioned GPUs in advance for potential workload bursts, which are underutilized for the most time. According to workload prediction and profiling, previous works can dynamically set the batch size and maximum waiting time to reduce the idle waiting (zhang2019mark, ; ali2020batch, ), but it can hardly be optimal. Some recent works (triton, ; pytorch2018pytorch, ; abadi2016tensorflow, ; zhai2022bytetransformer, ) bring in some operators supporting ragged batches to avoid extra paddings, but they can hardly support and perfectly optimize all operators and situations (DBLP:conf/mlsys/FegadeCGM22, ).
@@
In contrast, Student Parallelism trains a group of parallel students for GPUs other than a single student for weak edge devices, which can achieve the lowest model depth. It can even effectively reduce the model depth of large models like BERT-large and DeBERTa-xlarge, while the others fail. All students also have the same shallow architectures to be free of stragglers. Without over-provisioned GPUs, it can reduce the student number for higher throughput in case of workload bursts. Due to student parallelism, it also saves the costs of waiting queues and padding. Being orthogonal to quantization, it can quantize all students for further improvement.
@@ 3. Methodology
In this section, we first introduce the framework of Student Parallelism. Then we discuss how to offline train the equivalent group of parallel students for student parallelism. Finally, we describe how the multiple students run in parallel for online inference.
3.1. Framework Overview
We propose Student Parallelism to distill the original teacher model into a group of shallow student models for student parallelism, instead of a single student model. The core idea of Student Parallelism is to trade more parallel student models for less sequential layers to maintain accuracy. Specifically, Student Parallelism has the following two different phases to realize different goals:
  • Offline Training (§ 3.2): It aims to maintain the accuracy with the adaptive numbers of parallel students in low model depth. To be friendly to GPU acceleration, every student share the same shallow model architecture, resulting in weak model capability and homogeneity. To improve model capability and diversities, all students are sequentially trained on different sub-sampled subsets and virtually stacked via stacking distillation to gradually reduce the residual errors. By dynamically dropping some last-trained students, Student Parallelism conducts adaptive student pruning to improve the generalization of the remained students and minimize the accuracy loss. Then it can temporally reduce the student number to handle workload burst.
  • Online Inference (§ 3.3): it distributes the students of the same groups on the multiple GPUs to run in parallel for the single sample. Student Parallelism properly allocates the student group on GPUs and efficiently processes inference samples from online workload. And it replicates the student group into multiple copies to share GPUs with the support of MPS, so that Student Parallelism enables the immediate and concurrent inference of different individual samples without large data batches. If necessary, it employs the length-aware buffer to merge massive concurrent samples in similar lengths for higher throughput.
3.2. Offline Training
There are two challenges in training the equivalent group of parallel and shallow students for efficient GPU inference without affecting accuracy. One is how to maintain accuracy with those weak and homogeneous students having much fewer layers for low latency. The other is how to reduce student numbers with minimal accuracy loss for temporary high throughput. To overcome the two challenges, Student Parallelism employs the following two-stage training:
  1. (1)
    Sequential student training (line: 1-14): To generate a similar final pooled representation, all students are sequentially trained to gradually reduce the error between the teacher and all the students in boosting ensemble. To improve ensemble diversity and accuracy, it proposes different training losses, model initialization, and training datasets. All students are virtually stacked by the stacking distillation to improve the results.
  2. (2)
    Adaptive student pruning (line: 15-29): To temporarily improve throughput with minimal accuracy loss, Student Parallelism proposes adaptive student pruning to reduce the student number. By following the same way of sub-network training in NAS, it drops different numbers of last-added students to extract representations during the training. For every training data batch, it trains the output classifier upon different numbers of students to mimic the teacher’s prediction behavior.
3.2.1. Sequential Training of Student Models
Figure 3. The sequential training of virtually stacked students to gradually reduce the residual errors on the final representation between the teacher and all students.
Now we introduce how Student Parallelism trains the multiple parallel and shallow students to extract similar final representations with the original deep teacher. According to ensemble learning theory, different weak models need significant diversities to improve generalization and accuracy (zhou2021ensemble, ). Since all student models share the same shallow network for student parallelism, we propose the training strategy as described in Algorithm 1 to improve diversities from other aspects, including training loss, parameter initializations, and data subsets.
@@ 3.2.1. Sequential Training of Student Models
To achieve diversity, every student has different training loss functions applied to the final representation and intermediate representation (line 9) , as illustrate on Figure 3. All students are sequentially trained in the style of gradient boosting ensemble to gradually correct its own residual error on the final representation. Except that the first student directly mimics the final representation of the teacher, the $i-th$ student learns to correct the error of the ensemble of its own previous students. We keep adding new students until the overfitting happens according to the validation on spared test data (line 3). Formally, we define the boosting ensemble of the $M$ students having $N$ attention layers to generate the final representation as follows:
@@
In Equation 2, the $T(x)$ stands for the final representation output of the teacher model and $B^{(i-1)}(x)$ is the boosting ensemble of its all previous $i-1$ students. And only the i-th training student $S^{(i)}$ is trainable, while the other models like $T$ and $B^{(i-1)}(x)$ are fixed. After training the student $S^{(i)}$, we can set its multiplier $\alpha_{i}$ by line search to minimize the loss $\frac{1}{2}(T(x)-B^{(i)}(x))^{2}$, except that $\alpha_{0}$ is always 1. And we make the mathematical interpretation and formal proof on the better convergence of the final representations of all students based on the gradient boosting ensemble in the appendix.
@@
8:         outputs_t = TeacherModel(x)
9:         outputs_s = newStudent(x)
10:         Compute the losses as Equation 2 and Equation 3
11:         Update the newStudent
12:     end while
13:     studentList.append(copy(newStudent))
14:end while
15:studentList = studentList[:-1]
Despite that all students run independently, we propose stacking distillation to make every student virtually stacked on its own previous students. Any student learns to distill the ensemble of its all previous students into the representation of its $\lceil N/2\rceil$-th layer. Since stacking distillation makes the intermediate representation imitates all the previous students, the following layers can better correct the error by refining the intermediate representation, in the similar way of a deep BERT-like model. The loss $L_{\text{stack}}^{(i)}$ of stacking distillation for the i-th student is defined as follows:
@@
where $\lambda$ is the hyper-parameter to balance the final representation distillation and stacking distillation.
Besides the different training loss, Student Parallelism also brings in diversities from different perspectives. All students have different initialized parameters. The first student is initialized with a pre-trained language model with limited layers to improve the generalization, while any other student copies the parameters from the last trained student for the initialization (line 4). Additionally, every student model is trained on different sampled subsets of training data (line 5), except only the first student is trained on the whole training data. For any student, we pick up the top $a\%$ samples with the largest residual errors between the teacher and the existing students. And we randomly sample another $b\%$ of all the training data. And the subset of training data is augmented in the way suggested by previous works (jiao2019tinybert, ; hou2020dynabert, ).
3.2.2. Adaptive Pruning of Student Models

### v2

# Exploiting Student Parallelism for Efficient GPU Inference of BERT-like Models in Online Services

@@ Abstract.
Due to high accuracy, BERT-like models have been widely adopted by text mining and web searching. However, large BERT-like models suffer from inefficient online inference, facing the following two problems on GPUs: (1) their high accuracy relies on the large model depth, which linearly increases the sequential computation on GPUs; (2) stochastic and dynamic online workloads cause extra costs from batching and paddings.
@@ 1. Introduction
Due to the marvelous success in natural language understanding, BERT-like models (devlin2018bert, ; liu2019roberta, ; he2021deberta, ) have been the foundation model for discriminative tasks like text mining and web searching in online services. However, large BERT-like models pay the price of high inference time for the model capability. Besides high accuracy, a wide range of online services are time-sensitive, requiring efficient online inference. For example, news sentiment classification for automatic stock trading should be quicker than any other, because even a sub-millisecond delay can determine the profit (aquilina2022quantifying, ); for good user experiences, Retrieval Augmented Generation(RAG) (lewis2020retrieval, ) usually has a ten-millisecond-level time budget (yang2020model, ) for every query to search the relevant web pages as the reference of answer generation; and online writing assistants need to give immediate responses to text streams while users are typing (DBLP:conf/acl/KanekoMKSI20, ).
@@
Previous efficient inference works such as compact model design (sun2020mobilebert, ; xu2021bert, ), knowledge distillation (sanh2019distilbert, ; sun2019patient, ; jiao2019tinybert, ), and pruning (DBLP:conf/rep4nlp/GordonDA20, ; michel2019sixteen, ; hou2020dynabert, ) focus on reducing the model size for weak edge devices like cell phones. However, since deep learning essentially relies on model depth for high model capability (lecun2015deep, ), these methods prefer reducing model width to model depth (about half the layers remained) for better accuracy. Moreover, early existing methods (DBLP:conf/acl/XinTLYL20, ; DBLP:conf/acl/LiuZWZDJ20, ; zhou2020bert, ) still have some hard samples to go through all layers. Therefore, their deep architectures largely degrade their inference performance on parallel computing GPUs which are the de-facto device for online inference services. Figure 1(a) shows that model depth linearly increases the latency due to serial forward propagation. When GPU has enough parallel computation capability, it keeps the same latency and throughput for different model widths as shown in Figure 1(b). Therefore, these deep and thin models are not suitable for efficient GPU inference.
@@
Furthermore, existing works overlook the opportunity to optimize stochastic online workloads, resulting in the extra costs of batching and padding. Different from collected training data batches, online inference samples of different lengths arrive in stochastic and dynamic patterns. GPUs need large batches to achieve high throughput via data parallelism (kwon2023efficient, ), as Figure 1(c) shows that latency remains stable and throughput keeps increasing until the batch size is large enough. Therefore, it has to employ a waiting queue to batch enough data samples from stochastic online workloads (zhang2019mark, ; ali2020batch, ; DBLP:conf/usenix/Cui00WLZ0G22, ), and early-arrived samples have higher latency due to waiting. Text sequences have different lengths ranging from 10 to hundreds, but GPUs require all samples in one batch to have the regular length. Existing works commonly pad short samples into the maximum length to form the regular batch (devlin2018bert, ). As shown in Figure 1(d), latency can increase rapidly with sequence length due to $O(N^{2})$ complexity of self-attention (DBLP:conf/nips/ZaheerGDAAOPRWY20, ), so paddings also bring in significant wasted computations.
@@
Figure 1. Influences of different model and input factors on latency in red solid line) and throughput (green dashed line)
Therefore, we present Student Parallelism to explore student parallelism for efficient GPU inference of BERT-like models, especially for stochastic online workloads in practice. At its core, Student Parallelism distills the original deep model into an adaptive group of parallel and shallow student models, which are virtually stacked. Therefore, the sequential layer computations can be replaced with the equivalent parallel computations of students while maintaining accuracy. To achieve this, we propose a novel knowledge distillation method based on stacking distillation and explicit boosting ensemble(§ 3.1). Despite the parallel inference, every student is sequentially trained to gradually reduce the error between the teacher and the group of students. Every student distills all previous students into its intermediate layer by stacking distillation, so that it is virtually stacked on all the previous students. Then all students can make better corrections in the final layer with the estimated output of previous students. Consequently, all students can maintain accuracy with no more than one-quarter of the original layers (even only two layers sometimes, which is minimal for nonlinear classification), which significantly reduces the inference latency on GPUs. Beyond the distillation, it further conducts adaptive student pruning to improve generalization and reduce the student number. By dynamically dropping some last-trained students, Student Parallelism improves the generalization of the remained students, while pruning the redundant students causing over-fitting. More importantly, it realizes dynamically adjustable student numbers according to the changing workload. In case of workload bursts, it can temporarily reduce the student number to a small value(e.g., one or two) with minimal accuracy loss.
By leveraging the proposed student parallelism, we further make specialized designs to reduce waiting and padding from data parallelism. It allocates the students of the same group on different GPUs to better parallelize even the single sample. Then it replicates the student group into multiple copies to run concurrently and share the same set of GPUs. By employing the length-aware buffer, it enables immediate and concurrent inference of different samples to be free of batching large and regular data.
@@
As far as we know, Student Parallelism is the first to consider the real-world setting of efficient GPU inference on online workloads from the view of model design. Our technique contributions can be summarized as follows:
@@
    We leverage virtual stacking and boosting ensemble to distill the deep model into a group of parallel students in the lowest depth for low latency with comparable accuracy.
@@
    To handle online workloads, we can reduce the student number with minimal accuracy loss for workload bursts, and we also make specialized designs for student parallelism to reduce waiting and paddings.
@@
    Comprehensive experiments with real workload traces verify the effectiveness of our Student Parallelism, where it outperforms other baselines in terms of accuracy, latency, and throughput.
@@ 2. Related Work
  • Model Compression: these works only generate a single small model for weak edge devices, which still suffer from considerable model depth and high latency on GPUs. MobileBERT (sun2020mobilebert, ) and NAS-BERT (xu2021bert, ) make compact model design that can achieve the similar accuracy. Knowledge distillation transfers the knowledge from a large teacher (sanh2019distilbert, ; sun2019patient, ; jiao2019tinybert, ) or teacher ensemble (asif2020ensemble, ) to a small student model, so that the small student mimics the behavior of its teacher. Pruning can remove redundant weights, neurons, attention heads or tokens in well-trained models for less costs and better generalization (hou2020dynabert, ; malach2020proving, ; kim2022learned, ). Early exiting (DBLP:conf/acl/XinTLYL20, ; DBLP:conf/acl/LiuZWZDJ20, ; zhou2020bert, ) adds multiple classifiers on different layers as the multiple exits, so that it allows early exit from bottom layers for ”easy” samples. Quantization (kim2021bert, ; shen2020q, ; zafrir2019q8bert, ; tang2022mkq, ) replaces slow float operators with fast integer operators, which releases higher parallel computation capability. However, deep learning essentially relies more on model depth than model width for high model capability (lecun2015deep, ), so they still require a considerable large model depth to maintain the accuracy. For example, MobileBERT and NAS-BERT are still deep (12 or even 24 layers) but much thinner to avoid accuracy loss. 6-layer Distilbert can maintain 96.3% overall score of the 12-layer Bert teacher, but the 3-layer PKD-BERT can only get 88.5% (§ 4.2.1). Model pruning (hou2020dynabert, ) can only reduce about the half of layers at most to maintain good accuracy. In early exiting, the ”difficult” samples still have to go through all layers. Quantization not only needs hardware support in low-bit integers (zafrir2019q8bert, ) but also fails to reduce the model depth and computation complexity. It also relies on large batch sizes to increase parallelism degree and speed-up (e.g., the speedup is 2.4×2.4$\times$2.4 × and $\times3$ when the batch size is 1 and 8 respectively (kim2021bert, )).
@@
  • Model Ensemble: BERT stacks all attention layers with residual connections to enable the multi-path forward, which can be interpreted as the implicit boosting ensemble of all layers within the deep model (veit2016residual, ; huang2018learning, ). The Mixture-of-Experts (MoE) in attention heads (DBLP:conf/iclr/ShazeerMMDLHD17, ; DBLP:conf/naacl/ZuoZLHZC22, ) has a router to activate only one or two experts in every layer. It can keep constant inference costs no matter how wide the model is, but it cannot decrease the model depth. Bagging ensemble like Cocktail (DBLP:conf/nsdi/CrankshawWZFGS17, ; 276950, ) can select different relatively small models from the above-mentioned models to run in parallel for lower latency and high accuracy. However, they rely on the different model architectures for the ensemble diversity, resulting in the straggler problem on GPUs. SenseAI (wang2021sensai, ) divides the multi-class model into several pruned binary models, which still have a large model depth and irregular pruned architectures. It cannot be used in regression and binary classification tasks either. Progressive ensemble distillation (dennis2023progressive, ) can get closer to the teacher’s accuracy by adding more students with different networks. However, the maximum layer number of heterogeneous students remains to be large, and its accuracy decreases significantly if any students are dropped to speed up. Some Collaborative and Self-distillation (zhu2018knowledge, ; DBLP:conf/icann/ZainJZ22, ; sun2021collaborative, ) explore distilling multiple teachers and existing students into a new better student for higher accuracy, but they do not leverage multiple students together for lower model depth.
@@
  • Online Workloads Handling: Existing works in model design usually make the ideal assumption and ignore the fact that online workload is stochastic, so that online inference system has to pay extra costs of batching and padding in practice. They reserve some over-provisioned GPUs in advance for potential workload bursts, which are underutilized for the most time (zhang2019mark, ; DBLP:conf/usenix/Romero0YK21, ; 276950, ). According to workload prediction and profiling, previous system works can dynamically set the batch size and maximum waiting time to reduce the idle waiting (zhang2019mark, ; ali2020batch, ), but it can hardly be optimal. Some recent works (triton, ; pytorch2018pytorch, ; abadi2016tensorflow, ; zhai2022bytetransformer, ) bring in some operators supporting ragged batches to avoid extra paddings, but they can hardly support and perfectly optimize all operators and situations (DBLP:conf/mlsys/FegadeCGM22, ).
@@
In contrast, Student Parallelism trains a group of parallel students having the lowest model depth for GPUs, other than a single student for weak edge devices. Because of virtual stacking and boosting ensemble, it can even effectively reduce the model depth of large models like BERT-large and DeBERTa-xlarge, while the others fail. All students also have the same shallow architecture to be free of stragglers. Through adaptive student pruning, it can reduce the student number with minimal accuracy loss for higher throughput, being free of over-provisioned GPUs. Due to student parallelism, it also saves the costs of waiting queues and padding for the stochastic online workload. Being orthogonal to quantization, it can quantize all students for further improvement.
@@ 3. Methodology
In this section, we first show how to combine virtual stacking and boost ensemble to train the student group to generate final representations similar to deep teachers. Then we discuss how to apply adaptive student pruning in training the final classifier, which achieves better generalization and dynamical student numbers. Finally, we describe how the multiple students run in parallel for online inference.
3.1. Virtual Stacking and Boosting Ensemble
Figure 3. Virtually stacked students: the intermediate layer imitates all the previous students and the top layer reduces the residual error.
First of all, we recall that transformer models like BERT stack all the attention layers with residual connections. The residual connection adds the original input of the $i-th$ layer $x_{i}$ back to the result of its layer function $F_{i}(x_{i})$, so its output representation is $H_{i}=F_{i}(x_{i})+x_{i}$, namely the refinement of its input. Because there are numerous attention layers stacked with residual connections. Then the output of the n-th layer is expanded as $H_{n}=F_{n}(\sum_{i=1}^{n-1}H_{i}+x_{0})+\sum_{i=1}^{n-1}H_{i}+x_{0}$. Thus, residual connections enable the multi-path forward in the vision transformer, which behaves like the ensemble of networks in different depth (veit2016residual, ; huang2018learning, ).
Inspired by the implicit boosting ensemble within BERT, we propose to combine virtual stacking and boosting ensemble to trade more parallel student models for less sequential layers, while generating similar final representations of deep teachers. To be friendly to GPUs, all students share the same shallow model architecture, resulting in weak model capability and homogeneity. To improve model capability and diversities, every student is not only added to gradually reduce the error but also virtually stacked via stacking distillation to imitate the direct refinement of its own previous outputs, as illustrated in Figure 3. Overall, the training strategy as described in Algorithm 1 improves diversities from other aspects, including different training loss, different parameter initializations, and different data subsets.
@@ 3.1. Virtual Stacking and Boosting Ensemble
Boosting Ensemble. All students are sequentially trained in the style of boosting ensemble to gradually correct its own residual error on the final representation. Except that the first student directly mimics the final representation of the teacher, the $i-th$ student learns to correct the error of the ensemble of its own previous students. We keep adding new students until the overfitting happens on spared validation data (line 3 in Algorithm 1). Formally, we define the boosting ensemble of the $M$ students having $N$ attention layers to generate the final representation as follows:
@@
The $T(x)$ stands for the final representation output of the teacher model and $B^{(i-1)}(x)$ is the boosting ensemble of its all previous $i-1$ students. And only the i-th training student $S^{(i)}$ is trainable, while the other models like $T$ and $B^{(i-1)}(x)$ are fixed. After training the student $S^{(i)}$, we can set its multiplier $\alpha_{i}$ by line search to minimize the loss $\frac{1}{2}(T(x)-B^{(i)}(x))^{2}$, except that $\alpha_{0}$ is always 1.
@@
8:         $t\_rep_{N}$= TeacherModel(x)
9:         $prev\_rep=\sum_{j}^{k}\alpha_{j}studentList[j](batch)$
10:         $s\_rep_{N}$, $s\_rep_{\lceil N/2\rceil}$ = newStudent(x)
11:         Compute $L_{\text{boost}}$ as Equation 2 and $L_{\text{stack}}$ as Equation 3
12:         Update the newStudent
13:     end while
14:     studentList.append(newStudent)
15:end while
16:studentList = studentList[:-1]
Virtual Stacking. Despite that all students run independently, we propose stacking distillation to make every student virtually stacked on its own previous students. Any student learns to distill the ensemble of its all previous students into the representation of its $\lceil N/2\rceil$-th layer. Since stacking distillation makes the intermediate representation imitates all the previous students, the following layers can better correct the error by refining the intermediate representation, in a similar way to a deep BERT. The loss $L_{\text{stack}}^{(i)}$ of stacking distillation for the i-th student is defined as follows:
@@
where $\lambda$ is the hyper-parameter to balance the final representation distillation and stacking distillation. Since every student is trained under different virtual stacking terms $L_{\text{stack}}$, we improve model diversities by different training losses.
@@
Besides the different training losses, Student Parallelism also brings in diversities from different perspectives. All students have different initialized parameters. The first student is initialized with a pre-trained language model with limited layers to improve the generalization, while any other student copies the parameters from the last trained student for the initialization (line 4 in Algorithm 1). Additionally, every student model is trained on different sampled subsets of training data (line 5), except only the first student is trained on the whole training data. For any student, we pick up the top $a\%$ samples with the largest residual errors between the teacher and the existing students. And we randomly sample another $b\%$ of all the training data. The subset of training data is augmented in the way suggested by previous works (jiao2019tinybert, ; hou2020dynabert, ).
By following previous boosting ensemble theory (mason1999boosting, ; mason1999functional, ), we make mathematical proof on the better convergence of the final representations of all students based on the gradient boosting ensemble as follows:
Theorem 3.1.
The MSE in Student Parallelism has the lower bound 0, and it is a Lipschitz differentiable loss function (for any L ¿ 1, we always have $|\bigtriangledown L_{\text{boost}}(x_{1})-\bigtriangledown L_{\text{boost}}(x_{1})|<L|x_{1}-x_{2}|$). Let $F^{(0)},F^{(1)},...$ be the sequence of combined hypotheses generated by the Student Parallelism training algorithm, using small enough step-sizes $\alpha_{i}:=-\frac{\langle\bigtriangledown L_{\text{boost}}(T,B^{(i-1)}),S^{(i)}\rangle}{L|S^{(i)}|^{2}}$. Then our sequential training of students either halts on round T with $-\langle\bigtriangledown L_{\text{boost}}(T,F^{(i-1)}),S^{(i)}\rangle\leq 0$, or $L_{\text{boost}}(F^{(i)})$ converges to some finite value $L^{*}$, in which case
More proof details can be found in Appendix A.1.
3.2. Adaptive Student Pruning